OpenStudy (anonymous):

How can I convert y + 4 = (x-1) written standard form? I'm struggling with this partly because it doesn't seem to be in slope-intercept or point-slope form, because there is no slope/m value. I have no idea how to convert this. I have four choices for an answer and I've narrowed it down to two: -2x + y = -5 or -2x + y = -6 ^because they are the only choices in standard form. Can someone please help me convert this?

OpenStudy (freckles):

y+4=x-1 first write with x and y on the same side and all constant values on opposing side

OpenStudy (freckles):

not sure how you narrowed it down to a -2 in front of the x

OpenStudy (freckles):

unless y+4=(x-1) isn't what you meant to type

OpenStudy (anonymous):

I might have misunderstood it - the other options are y = 2x - 6 and y = 2x - 5. In class it was emphasized that it has to be in the Ax + By = C form, which the other two options didn't look to be

OpenStudy (anonymous):

I'm struggling to isolate the variables from the constants, and don't understand how many of the choices arrived at 2x anywhere in the equation.

OpenStudy (freckles):

so the equation really is y+4=(x-1) ?

OpenStudy (anonymous):

yes, it is

OpenStudy (freckles):

have you drop the ( ) because you don't need them y+4=x-1 since there is nothing acting upon them subtract y on both sides and add 1 to both sides and your equation will be in standard form

OpenStudy (anonymous):

thank you :)